The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                  http://www.astro.uiuc.edu/projects/mcnews/MCNews.php

No. 75 April 12th, 2004

Contents

14 Abstracts of Refereed Papers  2
1 Abstract of a Non-refereed Paper 11
1 Thesis Abstract  12
1 Meeting Announcement  13
1 Job Posting  14
Recent astro-ph Listings  15

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 75 of the Magellanic Clouds Newsletter. In this issue, we present 15 submitted abstracts, a thesis abstract from Yaël Nazé at Université de Liège, 2 announcements for an upcoming meetings, a new job listing from Geneva Observatory, and a collection of recent astro-ph listings of potential interest to Magellanic Clouds researchers. As always, this month’s issue is available from the MC News Website in a variety of formats (PDF, HTML, PostScript, and L^{A}TEX).

Best Regards, Bryan Dunne Editor, MC News

Abstracts of Refereed Papers

XMM-Newton observations of the giant HII region N 11 in the LMC

Y. Nazé (1), I.I. Antokhin (1,2), G. Rauw (1), Y.-H. Chu (3), E. Gosset (1), and J.-M. Vreux (1)

(1) Institut d’Astrophysique et de Géophysique, Université de Liège, Liège, Belgium

(2) Department of Physics and Astronomy, University of Glasgow, Glasgow, United Kingdom

(3) Astronomy Department, University of Illinois at Urbana-Champaign, Urbana, IL, USA

Using the sensitive XMM-Newton observatory, we have observed the giant HII region N 11 in the LMC for ∼30 ks. We have detected several large areas of soft diffuse X-ray emission along with 37 point sources. One of the most interesting results is the possible association of a faint X-ray source with BSDL 188, a small extended object of uncertain nature. The OB associations in the field-of-view (LH9, LH10 and LH13) are all detected with XMM-Newton, but they appear very different from one another. The diffuse soft X-ray emission associated with LH9 peaks near HD 32228, a dense cluster of massive stars. The combined emission of all individual massive stars of LH9 and of the superbubble they have created is not sufficient to explain the high level of emission observed: hidden SNRs, colliding-wind binaries and the numerous pre-main sequence stars of the cluster are most likely the cause of this discrepancy. The superbubble may also be leaking some hot gas in the ISM since faint, soft emission can be observed to the south of the cluster. The X-ray emission from LH10 consists of three pointlike sources and a soft extended emission of low intensity. The two brightest point sources are clearly associated with the fastest expanding bubbles blown by hot stars in the SW part of the cluster. The total X-ray emission from LH10 is rather soft, although it presents a higher temperature than the other soft emissions of the field. The discrepancy between the combined emission of the stars and the observed luminosity is here less severe than for LH9 and could be explained in terms of hot gas filling the wind-blown bubbles. On the other hand, the case of LH13 is different: it does not harbour any extended emission and its X-ray emission could most probably be explained by the Sk −66^{∘}41 cluster alone. Finally, our XMM-Newton observation included simultaneous observations with the OM camera that provide us with unique UV photometry of more than 6000 sources and enable the discovery of the UV emission from the SNR N11L.

Comments: Accepted by A&A WWW: http://vela.astro.ulg.ac.be/Preprints/P88/index.html or astro-ph/0402623 e-mail: naze@astro.ulg.ac.be

Tiny Scale Molecular Structures in the Magellanic Clouds. I. FUSE, HST , and VLT observations.

Martial-Kuassivi Andre et al.

Institut Astrophysique Paris

The objective of this series of two papers is to investigate small-scale molecular structures in the Magellanic Clouds (hereafter MCs). We report on the FUSE detections of the HD and CO molecules on the lines of sight towards three Large Magellanic stars: Sk −67D05, Sk −68D135, and Sk −69D246. HD is also detected for the first time on the lines of sight towards two Small Magellanic Cloud stars: AV 95 and Sk 159. While the HD and CO abundances are expected to be lower in the Large Magellanic Cloud where molecular fractions are a third of the Galactic value and where the photodissociation flux is up to thousands times larger, we report an average HD/H_{2} ratio of 1.4±0.5 ppm and CO/H_{2} ratio ranging from 0.8 to 2.7 ppm similar to the Galactic ones. We tentatively identify a deuterium reservoir (hereafter D–reservoir) towards the Small Magellanic Cloud, along the light path to AV 95. We derive a D/H ratio ranging from 1. 10^{−6} to 1.1 10^{−5}. Combining FUSE and HST/STIS data we also analyzed the H_{2}, Cl I, Cl II, Fe II, S II, C I, C I*, and C I** content, when available. High resolution VLT observations of Na I, Ca II, and Ca I were obtained in support of the lower resolution FUSE and STIS data for three targets in order to unravel the sightline velocity structures. These observations form the only such set of detections in the Magellanic Clouds to date and allow us to investigate in detail some of the physical properties of the intervening molecular gas. Our observation of the HD and CO molecules in the Magellanic Clouds is an argument for dense (n_{H} > 100 cm^{−3}) components. Furthermore, we demonstrate that these components are probably extremely small molecular clumps (possibly as small as 10^{2} pc) or filaments similar to the tiny–scale atomic structures (TSAS) recently observed in the halo of our Galaxy by Richter et al. (2003). Interestingly enough, for these five sightlines, we also detected molecular hydrogen originating in the Galactic disk. From these observations, we conclude that tiny–scale molecular filamentary structures are present in the disk of the Galaxy as well.

Comments: Paper accepted for publication in A&A Postscript/pdf version available in astro-ph e-mail: andre@iap.fr

A Chandra View of The Morphological And Spectral Evolution of Supernova Remnant 1987A

Sangwook Park(1), Svetozar A. Zhekov(2), David N. Burrows(1), Gordon P. Garmire(1), and Richard McCray(3)

(1) Penn State

(2) SRI, Bulgaria

(3) Colorado

We present an update on the results of our monitoring observations of the X-ray remnant of supernova (SN) 1987A with the Chandra X-Ray Observatory. As of 2002 December, we have performed a total of seven observations of SN 1987A, which allows us to monitor the details of the earliest stage of the supernova remnant evolution in X-rays. The high angular resolution images from the latest data reveal developments of new X-ray bright spots in the northwestern and the southwestern portions of the remnant as well as changes on the eastern side. The observed soft X-ray flux is increasing more rapidly than ever, and the latest 0.5−2 keV band flux (f_{X} ∼ 6 × 10^{−13} ergs cm^{−2} s^{−1}) is four times brighter than three years earlier when this monitoring began. The overall X-ray emission is primarily from the blast wave shock with kT ∼ 2.4 keV. As the blast wave approaches the dense circumstellar material, the contribution from the decelerated slow shock (kT ∼ 0.22 keV) to the observed X-ray emission is becoming significant. The increase of this slow shock contribution over the last two years is particularly noticeable in the western half of the remnant. These results indicate that the shock front is now reaching the main body of the inner circumstellar ring and that SN 1987A will be a complete ring with dramatic brightening in coming years. Based on the best-fit two-shock spectral model, we derive approximate densities of the X-ray-emitting regions (n_{e} ∼ 235 cm^{−3} for the fast shock and n_{e} ∼ 7500 cm^{−3} for the slow shock). There is no direct observational evidence to date for a neutron star associated with supernova remnant 1987A. We obtain an upper limit on the observed X-ray luminosity of any embedded point source (L_{X} ≤ 1.5 × 10^{34} ergs s^{−1}) in the 2−10 keV band. The X-ray remnant continues to expand linearly at a rate of 5129 km s^{−1}.

Comments: Accepted by ApJ (7/20/04 issue) e-mail: park@astro.psu.edu

Winds driven by super-star clusters: The self-consistent radiative solution

Sergey Silich, Guillermo Tenorio-Tagle, & Ary Rodriguez-Gonzalez

Instituto Nacional de Astrofisica Optica y Electronica, Puebla, Mexico

A self-consistent stationary solution for spherically symmetric winds driven by massive star clusters under the impact of radiative cooling is presented. We demonstrate that cooling may modify drastically the distribution of temperature if the rate of injected energy approaches a critical value. We also prove that the stationary wind solution does not exist whenever the energy radiated away at the star cluster center exceeds ∼30% of the energy deposition rate. Finally we thoroughly discuss the expected appearance of super-star cluster winds in the X-ray and visible line regimes. The three solutions here found: the quasi-adiabatic, the strongly radiative wind and the inhibited stationary solution, are then compared to the winds from Arches cluster, NGC 4303 central cluster and to the supernebula in NGC 5253.

Comments: Accepted for publication by The Astrophysical Journal e-mail: silich@inaoep.mx

Terminal velocities of luminous, early-type SMC stars

C. Evans, D. Lennon, C. Trundle, S. Heap, D. Lindler

Ultraviolet spectra from the Space Telescope Imaging Spectrograph (STIS) are used to determine terminal velocities for 11 O and B-type giants and supergiants in the Small Magellanic Cloud (SMC) from the Si IV and C IV resonance lines. Using archival data from observations with the Goddard High-Resolution Spectrograph and the International Ultraviolet Explorer telescope, terminal velocities are obtained for a further five B-type supergiants. We discuss the metallicity dependence of stellar terminal velocities for supergiants, finding no evidence for a significant scaling between Galactic and SMC metallicities for T_{eff} < 30,000 K, consistent with the predictions of radiation driven wind theory. A comparison of the v_{∞}/v_{esc} ratio between the SMC and Galactic samples, while consistent with the above statement, emphasizes that the uncertainties in the distances to galactic OB-type stars are a serious obstacle to a detailed comparison with theory. For the SMC sample there is considerable scatter in v_{∞}/v_{esc} at a given effective temperature, perhaps indicative of uncertainties in stellar masses.

Comments: Accepted by ApJ, to appear in May 20 edition. WWW: http://www.ing.iac.es/ cje/stis.ps.gz e-mail: cje@ing.iac.es

Puzzling wind properties of young massive stars in SMC-N81

F. Martins (1,2), D. Schaerer (2,1), D.J. Hillier (3), M. Heydari-Malayeri (4)

(1) LA2T-OMP, Toulouse, France

(2) Observatoire de Genève, Sayverny

(3) Department of Physics and Astronomy, University of Pittsburgh, Pittsburgh, PA, USA

(4) LERMA, Observatoire de Paris, Paris, France

We present a quantitative study of massive stars in the High Excitation Blob N81, a compact star forming region in the SMC. The stellar content was resolved by HST and STIS was used to obtain medium resolution spectra. The qualitative analysis of the stellar properties presented in Heydari-Malayeri et al. (2002) is extended using non-LTE spherically extended atmosphere models including line-blanketing computed with the code CMFGEN (Hillier & Miller 1998), and the wind properties are investigated. The main results are the following:

1) The SMC-N81 components are young (∼ 0–4 Myrs) O stars with effective temperatures compatible with medium to late subtypes and with luminosities lower than average Galactic O dwarfs, rendering them possible ZAMS candidates.

2) The winds are extremely weak: with values of the order 10^{−8}/10^{−9} M_{⊙} yr^{−1}, the mass loss rates are lower than o bserved so far for Galactic dwarfs. Only the recent study of SMC stars by Bouret et al. (2003) show the same trend. The modified wind momenta (Ṁ v_{∞} \sqrt{R}) are also 1 to 2 orders of magnitude lower than observed for Galactic stars. Both the mass loss rates and the modified wind momenta are lower than the predictions of the most recent hydrodynamical models.

The accuracy of the UV based mass loss rate determination, relying in particular on the predicted ionisation fractions, are carefully examined. We find that Ṁ could be underestimated by a factor of up to 10. Even in this unlikely case, the above conclusions remain valid qualitatively.

The reasons for such weak winds are investigated with special emphasis on the modified wind momenta:

1) There may be a break-down of the wind momentum - luminosity relation (WLR) for dwarf stars at low luminosity (log L/L_{⊙} ¡ 5.5). However, reasons for such a breakdown remain unknown.

2) The slope of the WLR may be steeper at low metallicity. This is predicted by the radiation driven wind theory, but the current hydrodynamical simulations do not show any change of the slope at SMC metallicity. Moreover, there are indications that some Galactic objects have wind momenta similar to those of the SMC stars.

3) Decoupling may take place in the atmosphere of the SMC-N81 stars, leading to multicomponent winds. However, various tests indicate that this is not likely to be the case.

The origin of the weakness of the wind observed in the SMC-N81 stars remains unknown. We suggest that this weakness may be linked with the youth of these stars and represents possibly the onset of stellar winds in recently formed massive stars.

Comments: Accepted by A&A WWW: http://fr.arxiv.org/abs/astro-ph/0403415 e-mail: fabrice.martins@obs.unige.ch

Massive Field Stars and the Stellar Clustering Law

M. S. Oey (1), N. L. King (1), and J. Wm. Parker (2)

(1) Lowell Observatory, Flagstaff, AZ, USA

(2) Department of Space Studies, Southwest Research Institute, Boulder, CO, USA

The distribution of N_{∗}, the number of OB stars per association or cluster, appears to follow a universal power-law form N_{∗}^{−2} in the local Universe. We evaluate the distribution of N_{∗} in the Small Magellanic Cloud using recent broadband optical and space-ultraviolet data, with special attention to the lowest values of N_{∗}. We find that the power-law distribution in N_{∗} continues smoothly down to N_{∗}=1. This strongly suggests that the formation of field massive stars is a continuous process with those in associations, and that the field stars do not originate from a different star formation mode. Our results are consistent with the model that field massive stars represent the most massive members in groups of smaller stars, as expected if the clustering law applies to much lower masses as is expected from the stellar initial mass function (IMF). These results are consistent with the simultaneous existence of a universal IMF and a universal clustering law. Jointly, these laws imply that the fraction of field OB stars typically ranges from about 35% to 7% for most astrophysical situations, with an inverse logarithmic dependence on the most populous cluster, and hence, on galaxy size and/or star formation rate. There are important consequences for global feedback effects in galaxies: field stars should therefore contribute proportionately to the volume of the warm ionized medium, and equal relative contributions by superbubbles of all sizes to the interstellar porosity are expected.

Comments: Accepted by AJ; available at astro-ph/0312051. WWW: http://www.lowell.edu/˜ oey; astro-ph/0312051 e-mail: sally.oey@lowell.edu

The Physical Properties and Effective Temperature Scale of O-type Stars as a Function of Metallicity. I. A Sample of 20 Stars in the

Magellanic Clouds

Philip Massey (1), Fabio Bresolin (2), Rolf P. Kudritzki (2), Joachim Puls (3), A. W. A. Pauldrach (3)

(1) Lowell Observatory

(2) Institute for Astronomy, University of Hawaii

(3) Universitats-Sternwarte Munchen

We have obtained HST and ground-based observations of a sample of 20 O-type stars in the LMC and SMC, including six of the hottest massive stars known (subtypes O2-3) in the R136 cluster. In general, these data include (a) the HST UV spectra in order to measure the terminal velocities of the stellar winds, (b) high signal-to-noise, blue-optical data where the primary temperature- and gravitysensitive photospheric lines are found, and (c) nebular-free Hα profiles, which provide the mass-loss rates. We find that the older (FOS) HST data of the R136 stars (which were obtained without the benefits of sky measurements) suffered from significant nebular emission, which would increase the derived mass-loss rates by factors of ∼ 3, all other factors being equal. We also find several stars in the SMC for which the N III λλ4634, 42 and He IIλ4686 emission “f” characteristics do not appear to follow the same pattern as in Galactic stars. Since He II emission is due to the stellar wind (which will be weaker in SMC for stars of the same luminosity), while N III emission is a complex NLTE effect affected mostly by temperature, it would not be surprising to find that these features do not correlate with each other or with luminosity in SMC stars in the same was as they do in Galactic stars, but theory does not provide a clean answer, and analysis of more stars (both SMC and Galactic) are needed to resolve this issue. The line-blanketed non-LTE atmosphere code FASTWIND was then used to determine the physical parameters of this sample of stars. We find good agreement between the synthetic line profiles for the hydrogen, He I, and He II lines in the majority of the stars we analyzed; the three exceptions show evidence of being incipiently resolved spectroscopic binaries or otherwise spectral composites. One such system is apparently an O3 V+O3 V eclipsing binary, and a follow-up radial velocity study is planned to obtain Keplerian masses. Although we did not use them to constrain the fits, good agreement is also found for the He I λ3187 and He II λ3203 lines in the near-UV, which we plan to exploit in future studies. Our effective temperatures are compared to those recently obtained by Repolust, Puls & Herrero for a sample of Galactic stars using the same techniques. We find that the Magellanic Cloud sample is 3,000-4,000^{∘}K hotter than their Galactic counterparts for the early through mid-O’s. These higher temperatures are the consequence of a decreased importance of wind emission, wind blanketing, and metal-line blanketing at lower metallicities.

Comments: Accepted by the Astrophysical Journal WWW: ftp://ftp.lowell.edu/pub/massey/haw.ps.gz e-mail: Phil.Massey@lowell.edu

An X-ray investigation of the NGC 346 field in the SMC (3): XMM-Newton data

Yaël Nazé (1), Jean Manfroid (1), Ian R. Stevens (2), Michael F. Corcoran (3), and Aaron Flores (4)

(1) Institut d’Astrophysique et de Géophysique, Université de Liège, Liège, Belgium

(2) School of Physics & Astronomy, University of Birmingham, Birmingham, United Kingdom

(3) Universities Space Research Association, Seabrook, MD, USA; and LHEA, GSFC, Greenbelt, MD, USA

(4) Universidad Autonoma del Carmen, Facultad de Ingenieria, Campeche, Mexico

We present new XMM-Newton results on the field around the NGC 346 star cluster in the SMC. This continues and extends previously published work on Chandra observations of the same field. The two XMM-Newton observations were obtained, respectively, six months before and six months after the previously published Chandra data. Of the 51 X-ray sources detected with XMM-Newton, 29 were already detected with Chandra. Comparing the properties of these X-ray sources in each of our three datasets has enabled us to investigate their variability on times scales of a year. Changes in the flux levels and/or spectral properties were observed for 21 of these sources. In addition, we discovered long-term variations in the X-ray properties of the peculiar system HD 5980, a luminous blue variable star, that is likely to be a colliding wind binary system, which displayed the largest luminosity during the first XMM-Newton observation.

Comments: Accepted by ApJ WWW: http://vela.astro.ulg.ac.be/Preprints/P89/index.html or astro-ph/0402629 e-mail: naze@astro.ulg.ac.be

A CNO Dichotomy among O2 Giant Spectra in the Magellanic Clouds

Nolan R. Walborn (1), Nidia I. Morrell (2), Ian D. Howarth (3), Paul A. Crowther (4), Daniel J. Lennon (5), Philip Massey (6), and Julia I. Arias (7)

(1) Space Telescope Science Institute, Baltimore, MD USA

(2) Las Campanas Observatory, The Carnegie Observatories, La Serena, Chile

(3) Department of Physics and Astronomy, University College London, London, UK

(4) Department of Physics and Astronomy, University of Sheffield, Sheffield, UK

(5) Isaac Newton Group, Canary Islands, Spain

(6) Lowell Observatory, Flagstaff, AZ, USA

(7) Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, La Plata, Argentina

From a survey of the 3400 Å region in the earliest O-type spectra, we have found that two of the four O2 giants observed in the Large Magellanic Cloud have the O IV lines there stronger than the N IV, while the other two have the opposite. A Small Magellanic Cloud counterpart also has the N IV stronger than O IV. Inspection of the blue spectra of these stars shows that the former pair have weaker N lines in all ionization states (III, IV, V) present as well as lines of C IV λ4658, while the latter three have stronger N lines and greater He/H. Space ultraviolet observations of two of the N-strong stars show N V wind profiles substantially stronger than those of C IV, while in the N-weak stars the C IV features are equal to or stronger than the N V. The N-strong stars are now reclassified as ON2 III(f*), newly defining that category. These characteristics strongly suggest a larger fraction of processed material in the atmospheres of the ON2 stars, which we confirm by modeling the optical spectra. In the context of current models, it is in turn implied that the ON2 stars are in a more advanced evolutionary state than the others, and/or that they had higher initial rotational velocities. The recent formulation of the effects of rotation on massive stellar evolution introduces an additional fundamental parameter, which the CNO abundances are in principle able to constrain. We present some illustrative comparisons with current Geneva evolutionary models for rotating massive stars. It is possible that these very hot, nitrogen-rich objects are products of homogeneous evolution. Our results will provide motivation for further physical modeling of the atmospheres and evolutionary histories of the most massive hot stars.

Comments: The paper is in press in the 2004 June 20 issue of ApJ, main journal. WWW: http://arXiv.org/abs/astro-ph/0403557 e-mail: walborn@stsci.edu

Surface Hydrogen Burning Modeling of Super Soft X-ray Binaries: They are Supernova Ia Progenitors

S. Starrfield (1), F. X. Timmes (2), W. R. Hix (3), E. M. Sion (4), W. M. Sparks (2), and S. J. Dwyer (1)

(1) Arizona State University,

(2) Los Alamos National Laboratory,

(3) Oak Ridge National Laboratory,

(4) Villanova University

Nova explosions occur on the white dwarf (WD) component of a Cataclysmic Variable stellar system which is accreting matter lost by a companion. A type Ia supernova explosion is thought to result when a WD, in a similar binary configuration, grows in mass to the Chandrasekhar Limit. Here, we present the first self-consistent calculations of accretion onto massive WDs which demonstrate that this evolution can occur. In this study, we have investigated the accretion of solar matter at a variety of mass accretion rates onto hot, luminous, WDs. In contrast to nova simulations where the WD has a low initial luminosity, the rate of mass transfer onto the WD is ≤ 10^{−9}M_{⊙} yr^{−1}, and material is periodically ejected at high velocities, these simulations do not eject material (or only a small fraction of the accreted material) and the WD grows in mass. Our simulations show that a hydrogen thermonuclear runaway cannot occur since hydrogen fuses immediately to helium in the surface layers, and we call this process Surface Hydrogen Burning (SHB). As the helium ash grows in mass, it gradually fuses both to carbon and more massive nuclei depending on the WD mass and mass accretion rate. If such a WD were to explode in a SN Ia event, therefore, it would show no hydrogen or helium in its spectra. Moreover, the surface conditions of our simulations agree with observations of some of the Super Soft X-ray Binary systems and, therefore, our results strengthen previous speculation that (some of) the Super Soft X-ray binaries (CAL 83 and CAL 87 for example) are progenitors of SN Ia explosions. Finally, we have achieved SHB for values of the mass accretion rate that span the observed values of the Cataclysmic Variables.

Comments: Submitted to ApJ Letters e-mail: starrfield@asu.edu

Stellar populations associated with the LMC Papillon Nebula

F. Meynadier (1), M. Heydari-Malayeri (1), L. Deharveng (2), V. Charmandaris (3 and 1), Th. Le Bertre (1), M.R. Rosa (4), D. Schaerer (5 and 6), H. Zinnecker (7)

(1) LERMA, Observatoire de Paris, Paris, France

(2) Observatoire de Marseille, France

(3) Cornell University, Astronomy Department, Ithaca, NY, USA

(4) Space Telescope European Coordinating Facility, European Southern Observatory, Garching bei München, Germany (5) Observatoire de Genève, Sauverny, Switzerland

(6) Laboratoire d’Astrophysique, Observatoire Midi-Pyrénées, Toulouse, France

(7) Astrophysikalisches Institut Potsdam, Potsdam, Germany

We study the Large Magellanic Cloud Papillon Nebula (N 159-5), a conspicuous High Excitation Blob (HEB) lying in the star forming complex N 159. Using JHK near-infrared photometry obtained at the ESO VLT with the ISAAC camera, we examine the stellar populations associated with the Papillon, tracing their history using stellar evolution models. Two populations are revealed: one composed of young, massive stars with an age ∼3 Myr, and a second consisting of older lower mass stars of age spreading between 1 and 10 Gyr. We analyze the properties of those populations and discuss their significance in the context of N 159. We also estimate that if the star at the center of the Papillon is single its initial mass is ∼50 solar masses, and it is affected by an extinction Av ∼7 mag.

Comments: Accepted by A&A WWW: http://arxiv.org/abs/astro-ph/0403626 e-mail: Frederic.Meynadier@obspm.fr

Low-Mass Pre-Main Sequence Stars in the Large Magellanic Cloud - III: Accretion Rates from HST-WFPC2 Observations

Martino Romaniello (1), Massimo Robberto (2,3) and Nino Panagia (2,3)

(1) European Southern Observatory, Garching bei München, Germany

(2) Space Telescope Science Institute, Baltimore, MD, USA

(3) Affiliated to the Astrophysics Division, Space Science Department of ESA

We have measured the present accretion rate of roughly 800 low-mass (∼ 1 − 1.4 M_{⊙}) pre-Main Sequence stars in the field of Supernova 1987A in the Large Magellanic Cloud (LMC, Z ≃ 0.3 Z_{⊙}). It is the first time that this fundamental parameter for star formation is determined for low-mass stars outside our Galaxy. The Balmer continuum emission used to derive the accretion rate positively correlates with the Hα excess. Both these phenomena are believed to originate from accretion from a circumstellar disk so that their simultaneous detection provides an important confirmation of the pre-Main Sequence nature of the Hα and UV excess objects, which are likely to be the LMC equivalent of Galactic Classical T Tauri stars. The stars with statistically significant excesses are measured to have accretion rates larger than ∼ 1.5 × 10^{−8}M_{⊙} yr^{−1} at an age of 12-16 Myrs. For comparison, the time scale for disk dissipation observed in the Galaxy is of the order of 6 Myrs. Moreover, the oldest Classical T Tauri star known in the Milky Way (TW Hydræ, with 10 Myrs of age) has a measured accretion rate of only 5 × 10^{−10} M_{⊙}/yr, i.e. 30 times less than what we measure for stars at a comparable age in the LMC. Our findings indicate that metallicity plays a major role in regulating the formation of low-mass stars.

Comments: Accepted for publication in the Astrophysical Journal (10 June 2004) WWW: http://arxiv.org/abs/astro-ph/0402459 e-mail: mromanie@eso.org

Photometry of Magellanic Cloud clusters with the Advanced Camera for Surveys - I. The old LMC clusters NGC 1928, 1939 and

Reticulum

A.D. Mackey and G.F. Gilmore

IoA Cambridge, UK

We present the results of photometric measurements from images of the LMC globular clusters NGC 1928, 1939 and Reticulum taken with the Advanced Camera for Surveys on the Hubble Space Telescope. Exposures through the F555W and F814W filters result in high accuracy colour-magnitude diagrams (CMDs) for these three clusters. This is the first time that CMDs for NGC 1928 and 1939 have been published. All three clusters possess CMDs with features indicating them to be > 10 Gyr old, including main sequence turn-offs at V ∼ 23 and well populated horizontal branches (HBs). We use the CMDs to obtain metallicity and reddening estimates for each cluster. NGC 1939 is a metal-poor cluster, with [Fe/H] = −2.10 ± 0.19, while NGC 1928 is significantly more metal-rich, with [Fe/H] = −1.27 ± 0.14. The abundance of Reticulum is intermediate between the two, with [Fe/H] = −1.66 ± 0.12 – a measurement which matches well with previous estimates. All three clusters are moderately reddened, with values ranging from E(V − I) = 0.07 ± 0.02 for Reticulum and E(V − I) = 0.08 ± 0.02 for NGC 1928, to E(V − I) = 0.16 ± 0.03 for NGC 1939. After correcting the CMDs for extinction we estimate the HB morphology of each cluster. NGC 1928 and 1939 possess HBs consisting almost exclusively of stars to the blue of the instability strip, with NGC 1928 in addition showing evidence for an extended blue HB. In contrast, Reticulum has an intermediate HB morphology, with stars across the instability strip. Using a variety of dating techniques we show that these three clusters are coeval with each other and the oldest Galactic and LMC globular clusters, to within ∼ 2 Gyr. The census of known old LMC globular clusters therefore now numbers 15 plus the unique, somewhat younger cluster ESO121-SC03. The NGC 1939 field contains another cluster in the line-of-sight, NGC 1938. A CMD for this object shows it to be less than ∼ 400 Myr old, and it is therefore unlikely to be physically associated with NGC 1939.

Comments: Accepted for publication in MNRAS e-mail: dmackey@ast.cam.ac.uk

Abstracts of Non-Refereed Papers

The Large Magellanic Cloud: Structure and Kinematics

Roeland P. van der Marel

Space Telescope Science Institute

I review our understanding of the structure and kinematics of the Large Magellanic Cloud (LMC), with a particular focus on recent results. This is an important topic, given the status of the LMC as a benchmark for studies of microlensing, tidal interactions, stellar populations, and the extragalactic distance scale. I address the observed morphology and kinematics of the LMC; the angles under which we view the LMC disk; its in-plane and vertical structure; the LMC self-lensing contribution to the total microlensing optical depth; the LMC orbit around the Milky Way; and the origin and interpretation of the Magellanic Stream. Our understanding of these topics is evolving rapidly, in particular due to the many large photometric and kinematic datasets that have become available in the last few years. It has now been established that: the LMC is considerably elongated in its disk plane; the LMC disk is thicker than previously believed; the LMC disk may have warps and twists; the LMC may have a pressure-supported halo; the inner regions of the LMC show unexpected complexities in their vertical structure; and precession and nutation of the LMC disk plane contribute measurably to the observed line-of-sight velocity field. However, many open questions remain and more work is needed before we can expect to converge on a fully coherent structural, dynamical and evolutionary picture that explains all observed features of the LMC.

Comments: Invited Review, 25 pages with 10 figures, to appear in ”The Local Group as an Astrophysical Laboratory”, Proceedings of the May 2003 STScI Symposium, M. Livio, ed., Cambridge, UK: Cambridge University Press. WWW: http://xxx.lanl.gov/abs/astro-ph/0404192 e-mail: marel@stsci.edu

Thesis Abstracts

Probing the Interactions of Early-Type Stars With Their Surroundings Through X-ray and Optical Observations

Yaël Nazé Advisor: Gregor Rauw

Institut d’Astrophysique et de Géophysique, Université de Liège, Liège, Belgium

This work is devoted to the study of the interactions of massive stars with their environment, either through their intense radiation fields or through their powerful stellar winds.

First, we have studied the close neighbourhood of the peculiar O star HD108, i.e. its stellar wind and associated emissions. We have notably uncovered the important variations of the emission lines seen in the visible domain. At the same time, in contrast, the X-ray emission remained stable. To better understand the X-rays emitted by the star, we have developed a program for modelling the ionisation of the stellar wind, that enabled us to determine the absorption of this ionised wind in the X-ray domain. Our X-ray observations have then permitted us to discard several models proposed to explain the nature of the system, and we have discussed the remaining possibilities in the light of the behaviour of the two other Of?p stars, the possible “twins” of HD108. The wind modelling program was also used to study the properties, in the X-ray range, of the Wolf-Rayet star WR40: combined to the observations taken by XMM-Newton, we have shown that the non-detection of X-rays from this star can be explained by the very large opacity of its wind. The study of the ionisation of the surrounding gas by Wolf-Rayet stars was extended through the analysis of HeII nebulae. For the first time, high quality images of the highly excited regions surrounding Wolf-Rayet stars were taken and analysed, enabling to derive the extreme UV flux of these stars, and thus their temperature, from ground-based observations.

In addition to the ionisation by massive stars, the impact of their stellar winds was also considered in this work, especially the wind-blown structures called “bubbles”. These bubbles arise from the action of the winds of massive stars, in isolation or in clusters, onto the interstellar medium. We first focused on the bubbles blown by isolated stars, and we started by the study of the high energy properties of the bubble blown by WR40. Thanks to XMM-Newton data, we showed that even the most recent theoretical models cannot account for the lack of X-ray emission of this bubble. To better understand the discrepancies between the theory and the observations of wind-blown bubbles, we then turned to the most simple structures, the ones blown by single main sequence massive stars. Several bubbles have been discovered in N11B, N180B, and N44, and their properties agree better with theoretical expectations than in the case of WR bubbles and superbubbles. Before our study, only a few interstellar bubbles were known: we showed that a simple morphological search was insufficient, and that kinematic studies were essential. In fact, the low-velocity expansion of these bubbles implies a weak compression of the gas, that can inhibit the morphological detection. Finally, we investigated the properties of the interaction of massive star clusters with their surroundings in the high-energy range. The clusters of N11 were observed in X-rays and UV thanks to the XMM-Newton satellite: their differences or similarities appear clearly at these energies, and we note that the diffuse X-ray emission always exceeded the expected level. Moreover, a unique dataset composed of Chandra and XMM-Newton observations enabled us to study the largest star formation region of the Small Magellanic Cloud, NGC346. Notably, the X-ray emissions from the NGC346 cluster itself and from the peculiar WR binary system HD5980 were discovered and analysed in details.

Our work underlines the complementarity between the visible, UV and X-ray domains, and the necessity to have observatories on the ground as well as in space. It also illustrates the important contribution of multiwavelength datasets to the study of massive stars. These observations indeed play a crucial role in the determination of the properties of early-type stars but they also provide critical tests for theoretical models (present or future).

e-mail: naze@astro.ulg.ac.be

Meeting Announcements

Astrophysics in the Far Ultraviolet

August 2-6, 2004 Victoria, BC, Canada

The Far Ultraviolet Spectroscopic Explorer continues to be a spectacularly successful Origins Mission, flown under the auspices of NASA, CNES (France) and the Canadian Space Agency. Operating in the 900-1200Åwavelength region, FUSE observations have resulted in a wealth of new insights into astronomical systems ranging from planets to the expanses of the intergalactic medium.

Hosted jointly by the Herzberg Institute of Astrophysics and the University of Victoria, a conference is to be held in Victoria, BC, Canada in 2004 to review the discoveries made with FUSE. The conference will highlight the role of far-ultraviolet data in multi-wavelength astronomical studies, and identify compelling scientific goals for future UV space missions.

Deadlines ∘ Abstract submission deadline: April 2, 2004 ∘ Registration deadline: April 30, 2004

For more information, contact the meeting organizers at: WWW: http://www.uvcs.uvic.ca/conf/FUSE/

WIDE-FIELD IMAGING FROM SPACE

May 16-18 2004 Lawrence Berkeley National Laboratory Berkeley, California, USA

The conference will address the scientific opportunities afforded by a wide-field optical and near-IR imager in space.

Conference Motivation:

Optical and near-infrared imaging from space would take a major step forward with the launch of wide-field instruments possessing sensitivities comparable to ACS or NICMOS on HST but covering hundreds of times the area. Such instruments would likely be carried on proposed missions to study the dark energy of the Universe. While the prime goal of such a mission would be the investigation of the equation of state of the universe, a significant fraction of the mission could be devoted to wide-field surveys proposed by community members, and could therefore open up entirely new classes of astronomical surveys. It is hoped that this conference may provide feedback that will help insure that proposed instruments will perform well both on the prime mission and on more general wide-field surveys. Talks will therefore span a large range of scientific issues. Complementary surveys from ground and space will also be discussed.

Conference Participation:

There will be about two dozen invited talks covering a wide range of astrophysics. However, contributions on scientific and technical issues relating to wide-field imaging from space are sought. The SOC expects there will be room for more than a dozen contributed talks. Substantial space and time will also be given over to poster contributions. Additionally, posters that are submitted in electronic as well as physical form will be available throughout the conference via a local wireless network.

Location:

The conference will be held on the grounds of the Lawrence Berkeley Laboratory, overlooking San Francisco bay. Hotel accommodation will be available at reduced rates at several hotels in Berkeley. A catered conference dinner will be held at the Laboratory. For other dinners the participants are invited to sample some of the excellent restaurants in the Berkeley area. Morning and afternoon snacks, as well as lunch on Sunday May 16 will be catered. Lunch and breakfast are available on the 17 and 18 May from the LBNL dining hall, which features a number of small eateries.

Deadlines ∘ March 31: Abstracts for oral presentations due ∘ April 12: Selection of contributed talks ∘ April 16: End early registration

For more information, contact the meeting organizers at: WWW: http://widefield.lbl.gov

Job Postings

Postdoctoral position on: Starbursts from the local Universe to high redshift

Postdoctoral position

The Geneva Observatory in Geneva, Switzerland, announces the availability of a research position at the postdoctoral level, open to applicants of all nationalities.

The succesfull candidate will work on a project aimed at studying starbursts nearby, at intermediate redshift (z∼2-3), or in the early Universe (z>7), involving multi-wavelength observations and/or theoretical modeling.The successfull applicant will in particular have access to ground-based observational data covering the optical to near-IR (including ESO/VLT data), and state-of-the-art modeling tools. He/she will mostly work in collaboration with Prof. Daniel Schaerer in Geneva.

The Geneva Observatory and the associated Laboratory of Astrophysics of the Swiss Federal Institute of Technology in Lausanne carry out observational, interpretative and theoretical research in the fields of extra-solar planets, stellar evolution, stellar physics, high energy astrophysics, galaxy evolution and dynamics, and observational cosmology.

The appointment will be for one to two years starting around October 1, 2004 (negotiable). Qualified candidates are encouraged to send their application including a CV and publication list,

description of research experience and interests, and contact information of three references preferably via email to the above address. All applications received by 1 May, 2004 will receive full consideration. Informal enquiries with Daniel Schaerer (daniel.schaerer@obs.unige.ch) are welcome.

Contact daniel.schaerer@obs.unige.ch

Related information also available from the URL http://obswww.unige.ch/sfr

Recent astro-ph Listings

astro-ph/0404231
Title: Quantitative studies of the Far-UV, UV and optical spectra of late O and
early B-type supergiants in the Magellanic Clouds
Authors: C. J. Evans, P. A. Crowther, A. W. Fullerton, D. J. Hillier
Comments: 47 pages, 15 figures (reduced size). Accepted for publication in ApJ.
A postscript version with full res. figs is available from this http URL
astro-ph/0404192
Title: The Large Magellanic Cloud: Structure and Kinematics
Authors: Roeland P. van der Marel (STScI)
Comments: Invited Review, 25 pages with 10 figures, to appear in "The Local
Group as an Astrophysical Laboratory", Proceedings of the May 2003 STScI
Symposium, M. Livio, ed., Cambridge, UK: Cambridge University Press. PDF file
with higher-resolution figures available at this http URL
astro-ph/0404098
Title: The distance to the LMC cluster Reticulum from the K-band
Period-Luminosity-Metallicity relation of RR Lyrae stars
Authors: M. Dall'ora (INAF-Oar, Univ. of Rome Tor Vergata), J. Storm (AIP),
G. Bono (INAF-Oar), V. Ripepi (INAF-Oac), M. Monelli (INAF-Oar, Univ. of
Rome Tor Vergata), V. Testa (INAF-Oar), G. Andreuzzi (INAF-Oar, TNG),
R. Buonanno (INAF-Oar, Univ. of Rome Tor Vergata), F. Caputo (INAF-Oar),
V. Castellani (INAF-Oar), C. E. Corsi (INAF-Oar), G. Marconi (ESO), M.
Marconi (INAF-Oac), L. Pulone (INAF-Oar), P. B. Stetson (DAO)
Comments: 14 pages, 2 figures, ApJ accepted. Full resolution figure 1 on request
(dallora@mporzio.astro.it)
astro-ph/0404063
Title: Tiny-Scale Molecular Structures in the Magellanic Clouds (Part 1)
Authors: Kuassivi-Andre, F. Lepetit, P. Sonnentrucker, R. Ferlet, E. Roueff, T.
Civeit, J-.M. Desert, S. Lacour, A. Vidal-Madjar
Comments: 34 pages, 10 tables, 12 figures, accepted for publication in A&A
astro-ph/0404017
Title: INTEGRAL observations of the Large Magellanic Cloud region
Authors: S.Mereghetti, D.Gotz, A.Paizis, A.Pellizzoni, S.Vercellone,
N.J.Westergaard, O.Vilhu, T.Belloni, R.Walter, T.Courvoisier, K.Ebisawa,
P.Kretschmar, L.Stella, J.-P.Swings, J.Knodlseder, A.Dean, A.Strong, P.Hakala,
A.Zdziarski
Comments: 4 pages, 7 figures. To be published in the Proceedings of the 5th
INTEGRAL Workshop: "The INTEGRAL Universe", February 16-20, 2004,
Munich

astro-ph/0403626
Title: Stellar populations associated with the LMC Papillon Nebula
Authors: F. Meynadier (1), M. Heydari-Malayeri (1), L. Deharveng (2), V.
Charmandaris (1 and 3), Th. Le Bertre (1), M.R. Rosa (4), D. Schaerer (5 and
6), H. Zinnecker (7) ((1) LERMA - Observatoire de Paris, France, (2)
Observatoire de Marseille, France, (3) Cornell University, Ithaca, USA, (4) ESO
- STECF, Munchen, Germany, (5) Observatoire de Geneve, Suisse, (6) OMP -
Laboratoire d'Astrophysique, Toulouse, France, (7) Astrophysikalisches
Institut, Potsdam, Germany)
Comments: Accepted in A&A 11 pages, 10 figures, 12 on-line material pages
(= 1 table) Some figures have been replaced by jpeg versions to reduce the
overall size
astro-ph/0403557
Title: A CNO Dichotomy among O2 Giant Spectra in the Magellanic Clouds
Authors: Nolan R. Walborn, Nidia I. Morrell, Ian D. Howarth, Paul A. Crowther,
Daniel J. Lennon, Philip Massey, Julia I. Arias
Comments: 15 pages, 8 figures, 1 table
astro-ph/0403415
Title: Puzzling wind properties of young massive stars in SMC-N81
Authors: F. Martins, D. Schaerer, D.J. Hillier, M. Heydari-Malayeri
Comments: 23 pages, 22 figures, A&A accepted
astro-ph/0403410
Title: Probing the RGB-phase transition: Near-IR photometry of six
intermediate age LMC clusters
Authors: Francesco R. Ferraro, Livia Origlia, Vincenzo Testa, Claudia
Maraston
Comments: ApJ, in press
astro-ph/0403343
Title: Raising the Dead: Clues to Type Ia Supernova Physics from the Remnant
0509-67.5
Authors: Jessica S. Warren, John P. Hughes (Rutgers University)
Comments: 34 pages, 9 figures (1 color), accepted to ApJ (10 June 2004 issue)
astro-ph/0403331
Title: Core Formation by a Population of Massive Remnants
Authors: David Merritt, Slawomir Piatek, Simon Portegies Zwart, Marc
Hemsendorf
Comments: 5 pages
astro-ph/0403069
Title: Two Years of Observations of the X-ray Pulsar SMC X-1 with the
ART-P Telescope onboard the Granat Observatory
Authors: A.A.Lutovinov (1,2), S.S.Tsygankov (1), S.A.Grebenev (1),
M.N.Pavlinsky (1), R.A.Sunyaev (1,2) (1 - Space Research Institute, Moscow,
2 - MPA, Garching)
Comments: 16 pages, 5 figures
Journal-ref: Astronomy Letters, 2004, V.30, pp.50-57
astro-ph/0402633
Title: The Physical Properties and Effective Temperature Scale of O-type Stars
as a Function of Metallicity. I. A Sample of 20 Stars in the Magellanic Clouds
Authors: Philip Massey, Fabio Bresolin, Rolf P. Kudritzki, Joachim Puls, A. W.
A. Pauldrach
Comments: Accepted for publication in the Astrophysical Journal. A postscript
version with the figures embedded can be found at this ftp URL

astro-ph/0402629
Title: An X-ray investigation of the NGC 346 field in the SMC (3):
XMM-Newton data
Authors: Yal Naz (1), Jean Manfroid (1), Ian R. Stevens (2), Michael F.
Corcoran (3), Aaron Flores (4) ((1) Institut d'Astrophysique et de
Geophysique, Liege, Belgium; (2) University of Birmingham UK ; (3) GSFC,
USA; (4) Universidad Autonoma del Carmen, Mexico)
Comments: 19 pages, 5 figures (in gif), accepted by ApJ, also available from this
http URL
astro-ph/0402623
Title: XMM-Newton observations of the giant HII region N11 in the LMC
Authors: Y. Naze (1), I.I. Antokhin (1,2), G. Rauw (1), Y.-H. Chu (3), J.-M.
Vreux (1) ((1) Institut d'Astrophysique et de Geophysique, Liege, Belgium; (2)
University of Glasgow, UK ; (3) University of Illinois at Urbana-Champaign,
USA)
Comments: 13 pages, 16 figures (10 in gif or jpg), accepted by A&A, also
available from this http URL
astro-ph/0402621
Title: Two long-period X-ray pulsars detected in the SMC field around XTE
J0055-727
Authors: F. Haberl, W. Pietsch, N. Schartel, P. Rodriguez, R.H.D. Corbet
Comments: Submitted to A&A Letters
astro-ph/0402559
Title: Symbiotic Stars in the Magellanic Clouds
Authors: Joanna Mikolajewska
Comments: 2 pages, 4 figues To appear in IAU Coll. No. 194: Compact Stars in
the Galaxy and Beyond, RevMexAA(SC)
astro-ph/0402459
Title: Low-Mass Pre-Main Sequence Stars in the Large Magellanic Cloud -
III: Accretion Rates from HST-WFPC2 Observations
Authors: Martino Romaniello, Massimo Robberto, Nino Panagia
Comments: Accepted for publication in the Astrophysical Journal (10 June 2004),
28 pages, 9 figures. Typo corrected in the abstract on 21 February 2004
astro-ph/0402424
Title: New Period-Luminosity and Period-Color Relations of Classical
Cepheids: II. Cepheids in LMC
Authors: A. Sandage (1), G.A. Tammann (2), B. Reindl (2) ((1) Obs. Carnegie
Inst. Washington, (2) Astr. Inst. Univ. Basel)
Comments: 50 pages, 20 figures, 5 tables. Submitted to Astronomy and
Astrophysics
astro-ph/0402187
Title: High-resolution N-body Simulations of Galactic Cannibalism: The
Magellanic Stream
Authors: Tim W. Connors (1), Daisuke Kawata (1), Sarah T. Maddison (1),
Brad K. Gibson (1) ((1) Swinburne)
Comments: 6 pages, 4 figures, LaTeX (gcdv.sty). Refereed contribution to the
5th Galactic Chemodynamics conference held in Swinburne, July 2003. Accepted
for publication in PASA. Version with high resolution figures available at this
http URL